Suppose a van de Graaff generator builds a negative staticcharge, and a grounded conductor is placed near enough to it sothat a 7μC of negative charge arcs to the conductor.Estimate the number of electrons involved.

4.38 × 1013

112

7 × 1020

1
Answer #1
Charge Q = 7 * 10 ^ -6 C
we know Q = N e
from this no.of electrons involved N = Q / e
                                                        =[ 7 * 10 ^ -6 C ] / [ 1.6 * 10 ^ -19 C ]
                                                      = 4.375 * 10 ^ 13
                                                       ~ 4.38 * 10 ^ 13